An object displaces 2.3×10-³ m-³ of water when totally immersed into it. What is the bouyant force exerted by water? What is the density of the object in its mass is 26 kg? Of what material is the object was made of?


1
Expert's answer
2022-03-31T13:45:01-0400

The buoyant force is the weight of the water displaced, which is

(water density)(volume)*g

= 2.3*10^-3 m^3*1.00*10^3 kg/m^3*9.8 m/s^2 = 22.5 Newtons


The density is (26,000 g)/(2.3^10^3) cm^3 = 11.3 g/cm^3


That happens to be the density of the metal lead. There are probably some alloys that have the same density, but they are not commonly used.